# Composing Agents to Minimize Worst-case Risk

### v1

@@ 3 Finding the path minimizing the worst-case risk
6$\triangleright$ Map to store samples along best partial path
7:$\textnormal{{VaR}}[s,\_]\leftarrow-\infty;\texttt{best\_paths}[s,\_]\leftarrow[0,0]$triangleright Base case: no risk at source
@@
12$\triangleright$ All predecessors have been processed due to topological order
@@
14:     for $v^{\prime}\in\texttt{pred}$ and $\alpha^{\prime}\in B_{\leq\bar{\alpha}}$ $\triangleright$ $B_{\leq\bar{\alpha}}\coloneqq\{q\in B:q\leq\bar{\alpha}\}$
@@
18$\triangleright$ Sample trace-output pairs along edge
@@
[…]
21:         $\texttt{pathVaR}\leftarrow\max(\texttt{VaR}[v^{\prime},\alpha^{\prime}],\texttt{edgeVaR})$triangleright Worst-case risk along the path
@@
23:              $\textnormal{{VaR}}[v,\bar{\alpha}]\leftarrow\texttt{pathVaR}$triangleright Update best estimate
[…]
@@
29:return $\textnormal{{VaR}}[t,\alpha],\texttt{best\_paths}[t,\alpha]$triangleright Path with minimum VaR estimate
Towards optimizing the WoRM objective over an agent graph $G$, we remark that a simple procedure to find the optimal path is to estimate the $\textnormal{VaR}_{\alpha}$ along each path and pick the path that minimizes it. We can estimate the $\textnormal{VaR}_{\alpha}$ by computing the empirical $(1-\alpha)$-quantile on samples drawn from the loss along a path $L_{p}(Z_{p})$. While this is asymptotically optimal, it can be inefficient when the graph has an exponential number of paths as a function of the number of vertices.
To circumvent this, we present an efficient algorithm that only depends polynomially on the number of vertices in the agent graph and is also asymptotically near-optimal under certain assumptions. BucketedVaR is a dynamic programming algorithm that processes the vertices in the agent graph in the topological order and incrementally builds a near-optimal path.
To understand how it works, we first present how we can estimate the value-at-risk of the loss distribution of a single path incrementally. To this end, consider a path in an agent graph $p=v_{1}\xrightarrow{e_{1}}\dots\xrightarrow{e_{m}}v_{m+1}$ with $v_{1}$ being the source vertex and $v_{m+1}$ the target vertex. Let $(Z_{1},\dots,Z_{m})\sim\mathcal{D}(T_{e_{1}}\times\cdots\times T_{e_{m}})$ be a sequence of random variables representing the sequences of traces of the agents along the path. Then define $R_{1},\dots,R_{m}$ to be real-valued random variables representing the losses incurred by the agents along each edge. Formally, we can write $R_{1}\coloneqq L_{e_{1}}(Z_{1}),\dots,R_{m}\coloneqq L_{e_{m}}(Z_{m})$. If $L_{p}$ is the composed loss function along the path as defined in Equation 2, we can rewrite $L_{p}(Z)$ as $\max(R_{1},\dots,R_{m})$. A key observation that enables the incremental estimation of $\textnormal{VaR}_{\alpha}[L_{p}(Z)]$ (or equivalently, the $(1-\alpha)$-quantile of $L_{p}(Z)$) is that we can estimate $\textnormal{VaR}_{\alpha_{i}}[R_{i}]$ of each loss variable individually with $\alpha_{i}$ chosen such that $\sum\alpha_{i}=\alpha$. Following this, we can apply the union bound to obtain
@@
This bound is justified in the proof of Theorem 1. This method allows us to estimate the value-at-risk of each edge loss variable independently and then combine the estimates at the end by computing their maximum. However, we need to optimize the allocation of the risk budgets $\alpha_{i}$ along each edge. Intuitively, allocating a higher risk budget to higher edge losses would result in a lower value-at-risk for that edge since it would correspond to a lower quantile.
BucketedVaR optimizes this by searching for risk budget allocations in a discretized set of buckets $B=\{0,\frac{\alpha}{d},2\frac{\alpha}{d},\dots,\alpha\}$ for a given discretization factor $d\in\mathbb{N}_{>0}$. At each vertex $v$ and corresponding to each bucket $\bar{\alpha}\in B$, the algorithm inductively builds an optimal path that minimizes the upper bound on the value-at-risk along the partial path at a risk budget of $\bar{\alpha}$. It achieves this by considering each predecessor $v^{\prime}$ of $v$ and each allocation of budget $\alpha^{\prime}\leq\bar{\alpha}$ until $v^{\prime}$, for which we have inductively constructed the optimal path, and then estimates the value-at-risk by allocating a budget of $\bar{\alpha}-\alpha^{\prime}$ to the loss along the edge $v^{\prime}\to v$. The optimal path for $(v,\bar{\alpha})$ is then obtained by extending the optimal path to $(v^{\prime},\alpha^{\prime})$ that minimizes the value-at-risk estimate. A detailed proof of soundness and run-time analysis of the algorithm can be found in the proof of Theorem 1
@@
Consider $G=(V,E,X,T,F,L,s,t,\mathcal{D}_{s})$ an agent graph, $d\in\mathbb{N}_{>0}$ the number of buckets, $n\in\mathbb{N}_{>0}$ the sample size, and $\alpha\in(0,1)$ the risk budget. Then, let $q\in\mathbb{R}$ and $p\in V^{*}$ be the value-at-risk estimate and path returned by BucketedVaR$(G,d,n,\alpha)$. Then, for all $\delta\in(0,1)$, with probability at least $1-\delta$,
@@
Furthermore, the time complexity of BucketedVaR$(G,d,n,\alpha)$ is $O(n(d+1)^{2}\lvert V\rvert^{2})$ assuming that sampling from an agents’ output distributions and the initial distribution $\mathcal{D}_{s}$ incur constant costs.
@@
Observe that since $\gamma\to 0$ as $n\to\infty$, the theorem tells us that the estimated value-at-risk $q$ is at least as large as the true value-at-risk of the loss distribution of the path $p$, which corresponds to $\textnormal{quantile}(L_{p}(Z_{p}),1-\alpha)$. We can show this using a union bound argument for the budget allocations found by the algorithm along with the fact that the empirical CDF converges uniformly to the true CDF by the DKW inequality [8, 23]. The time complexity is easily deduced by looking at the pseudocode of BucketedVaR. See appendix for the complete proof. ∎
Theorem 1 only justifies the validity of the union bounding method for estimating the value-at-risk but we have not discussed the optimality of BucketedVaR. Here, we can only hope to achieve tight estimates of the value-at-risk when the loss variables along every path are independent or are only loosely correlated. Under these independence assumptions, Theorem 2 tells us that asymptotically, the quantile estimated by the BucketedVaR algorithm is at most the $(1-\alpha+\alpha^{2}/2)$-quantile of the losses along the optimal path. In other words, the path found by our algorithm is suboptimal by an additive factor of $\alpha^{2}/2$ with respect to the quantiles of the optimal path loss.
@@
This is again a consequence of the uniform convergence of the empirical CDF along with the fact that the independence of the loss variables allows us to deduce an upper bound. See appendix for the complete proof. ∎
@@ 4 Experimental evaluation
We implement  [^1] The implementation can be found at https://github.com/guruprerana/worm the BucketedVaR algorithm for worst-case risk minimization and evaluate its performance on a suite of discrete and continuous control benchmarks that require composing agents trained with reinforcement learning (RL) to complete long-horizon objectives. In all the benchmarks that we consider, the agent graphs $G=(V,E,X,T,F,L,s,t,\mathcal{D}_{s})$ share the following structure: the domain associated with each vertex $X_{v}$ is the state space of the control environment $\Sigma$ and the agents along each edge are control policies $\pi_{e}:\Sigma\to\mathfrak{D}(\Sigma^{*}\times\Sigma)$ that take an initial state and produce a distribution over trajectories and final states. These final states then become the initial states for the next policy. Lastly, $\mathcal{D}_{s}$ represents the distribution over the initial states of the environment.
@@
Maximizing safety during navigation and manipulation. We test our algorithm on the 16-Rooms and Fetch environments from Jothimurugan et al. [17] who present a compositional RL framework to complete long-horizon objectives in continuous environments. An objective is specified as a task graph which is, in essence, the same as an agent graph. In the task graph, every edge represents a reach-avoid subtask: reaching a target region while avoiding a set of dangerous states. Each reach-avoid subtask specifies a real-valued reward function that takes in a trajectory of execution of the policy and evaluates to at least zero if the trajectory successfully completes the subtask. Intuitively, the reward captures a distance metric between trajectories and reach-avoid sets. A task graph can be naturally mapped to an agent graph: the loss function is defined as the negative of the reward function.
@@
The 16-Rooms environment looks similar to the MouseNav environment and consists of 16 rooms arranged in a $4\times 4$ grid with doors between adjacent rooms. RL policies are trained to control a point mass to complete reach-avoid tasks. The agent graph has 13 vertices and 16 paths of length 8 each. The second Fetch environment [26] is taken from Gymnasium-Robotics and involves controlling a 7-DoF robotic arm with a gripper attached at the end to pick up and move objects. In this environment, we consider the task graph for the PickAndPlaceChoice objective that directs the gripper to move near the object, grip it, pick it up and move along one of two trajectories. The graph has 7 vertices and 2 paths of length 5. Additionally, we adapt the Rooms environment to implement the MouseNav environment from Figure 1.
(a) Top-view map with agent graph
(b) First-person view when carrying box
Figure 2: BoxRelay environment
@@
Minimizing resource consumption. We use the Miniworld [5] framework to implement the BoxRelay environment of which the top-view map is shown in Figure 2(a). Observations in this environment are given as first-person view RGB images as seen in Figure 2(b). The objective: we have to move the player (red triangle) to the first box (marked 1) and pick it up. The player can then choose to go to either of the two boxes in the two middle rooms (3 or 7), place the initial box, pick up the next box, and finally drop it at the location marked in the right room with a star. Furthermore, after picking up the box at 7, the player can choose to exit through one of the two exits 8 or 9. We add the constraint that the player has limited battery power but gets recharged after dropping the first box and before picking up the second box. So we would like to minimize the amount of time that either box is held by the player. Importantly, since the initial positions of the player, the boxes, and the final target are all chosen uniformly at random within each room, this is a non-trivial optimization problem.
@@
We can model the BoxRelay task as an agent graph as follows: the graph is as shown in Figure 2(a) with the numbers being the vertices. The edges represent the agents which are RL policies trained to navigate the player to pick up and place objects at the desired locations. The loss functions evaluate to the number of time steps that a box is carried before being put back down and thus quantifies the resource consumption. Here, the WoRM objective corresponds to minimizing the maximum time that any box is carried by the player.
@@
Equipped with these benchmark environments, we conduct three sets of experiments to evaluate the effectiveness of the approximation obtained from the BucketedVaR algorithm. We measure the accuracy of the approximation by estimating the empirical coverage on $10^{4}$ fresh samples of losses along the optimal path. Specifically, if $c\in\mathbb{R}$ is the approximate $\textnormal{VaR}_{\alpha}$ returned by the BucketedVaR algorithm, we draw samples of losses along the optimal path and compute the fraction of these which are below $c$ giving us the empirical coverage of $c$. We say that the approximation is tight if the empirical coverage is close to the desired quantile $1-\alpha$.
Table 1: Comparison of BucketedVaR with the optimal baseline algorithm on $10^{4}$ samples. We prefer empirical coverages that are close to $1-\alpha$ where $\alpha$ is the risk level. 95% Clopper-Pearson CIs are specified for the empirical coverages computed on a fresh set of $10^{4}$ samples.
Benchmark
Risk
level
Buckets
BucketedVaR
estimate
Baseline
estimate
BucketedVaR
coverage (%)
Baseline
coverage (%)
MouseNav 0.2 5 -0.339 -0.339 79.56[78.75, 80.34] 79.56[78.75, 80.34]
0.1 5 -0.330 -0.330 89.77[89.15, 90.35] 89.77[89.15, 90.35]
0.05 5 -0.328 -0.328 95.01[94.56, 95.42] 95.01[94.56, 95.42]
0.1 100 -0.0135 -0.0135 91.24[90.66, 91.78] 91.24[90.66, 91.78]
0.05 100 -0.0084 -0.0084 95.45[95.02, 95.85] 95.45[95.02, 95.85]
Fetch 0.2 30 0.0314 0.0250 80.45[79.65, 81.22] 79.58[78.77, 80.36]
0.1 30 0.3483 0.3393 90.92[90.33, 91.47] 90.53[89.93, 91.09]
0.05 30 0.3914 0.3881 95.40[94.97, 95.80] 95.28[94.84, 95.68]
BoxRelay 0.2 50 151 148 81.79[81.01, 82.54] 80.56[79.77, 81.33]
0.1 50 176 175 90.90[90.31, 91.45] 90.70[90.11, 91.26]
0.05 50 210 208 95.46[95.03, 95.85] 95.34[94.90, 95.74]
@@
Figure 3: Empirical coverage of the approximation of $\textnormal{VaR}_{0.1}$ from BucketedVaR with varying parameters. $95\%$ Clopper-Pearson CIs for the coverage computed on $10^{4}$ samples are also plotted.
Comparison with optimal baseline algorithm. We define a baseline algorithm that explicitly esimates $\textnormal{VaR}_{\alpha}$ at the desired risk level $\alpha$ along each path of the agent graph. While this algorithm is asymptotically optimal, it is inefficient as discussed in the previous section. Table 1 shows the VaR of the paths found by both BucketedVaR and the baseline algorithm on all of the benchmarks at 3 different risk budgets $\alpha=0.2,0.1,0.05$. The number of samples for both algorithms is set to $10^{4}$ and the number of buckets is chosen based on the number of agents to allow sufficient granularity in the allocation of the risk budget. For example, in the 16-Rooms benchmark since there are 8 agents along each path, we choose a higher number of buckets.
We find that BucketedVaR succeeds in finding the optimal path across all the benchmarks along with tight estimates of $\textnormal{VaR}_{\alpha}$. It also finds non-trivial allocations the risk budget for the agents: in the 16-Rooms benchmark, the budget allocations for the 8 agents along the path to estimate $\textnormal{VaR}_{0.1}$ are $16\bar{\alpha},0\bar{\alpha},10\bar{\alpha},23\bar{\alpha},19\bar{\alpha},11\bar{\alpha},7\bar{\alpha}$, and $14\bar{\alpha}$ for $\bar{\alpha}=0.1/100$.
Improving approximation with increasing sample size and number of buckets. Increasing the sample size improves the approximation because we have more accurate quantile estimates and similarly a higher number of buckets results in more granular risk budget allocations. This is confirmed by the two experiments we run: in the first one, plotted in Figure 3(a), we maintain the number of buckets at 100, and vary the number of samples from 500 to $10^{4}$ with $\alpha=0.1$. The empirical coverage quickly stabilizes around 0.91 across benchmarks. The second experiment maintains the sample size at $10^{4}$ and varies the number of buckets from 5 to 100 with $\alpha=0.1$. Figure 3(b) again shows a similar behavior. We note that for the 16-Rooms benchmark, a low number of buckets gives bad empirical coverage since the number of agents along each path is higher.
Scaling number of agents. From our theoretical result, we expect BucketedVaR to continue producing tight estimations of the VaR when we increase the number of agents along the path, provided that the condition on the independence of losses holds and the sample size is large enough. We test this hypothesis by simulating a long path of agents using the 16-Rooms benchmark. We consider the optimal path in this benchmark which has a sequence of 8 agents and construct a new agent graph with $8k$ agents for $k\in\{1,2,3,4,5\}$. We can sample a sequence of trajectories from this path by taking $k$ trajectories of the original 8 agents. The losses are then computed using the original loss functions. Since these are independently drawn trajectories, the assumption of independence holds. We estimate the $\textnormal{VaR}_{\alpha}$ of this path of agents using BucketedVaR for $10^{4}$ samples, $\alpha=0.1$, and $100$ buckets. The results are plotted in Figure 3(c) in which we observe that the approximation continues to remain tight with an increasing number of agents.
Takeaways. Through our benchmarks, we have illustrated that we can use our framework to model risk associated with agent compositions like safety and resource consumption. Then, as seen in Table 1, we showed that the BucketedVaR algorithm produces estimates of the value-at-risk that are tight with respect to the optimal estimate and also achieves empirical coverage very close to the desired level of $1-\alpha$. Furthermore, we are able to confirm our theoretical result that the approximation improves with growing sample size and granularity of the risk budget division given by the number of buckets. We also observe that it scales to longer paths with more agents.

### v2

@@ 3 Finding the path minimizes the value-at-risk
[…]
@@
[…]
@@
20:   $\texttt{edgeVaR}\leftarrow$ $\textnormal{quantile}(\texttt{losses},1-(\bar{\alpha}-\alpha^{\prime}))$ $\triangleright$ Empirical $(1-(\bar{\alpha}-\alpha^{\prime}))$-quantile
21:   $\texttt{pathVaR}\leftarrow\max(\texttt{VaR}[v^{\prime},\alpha^{\prime}],\texttt{edgeVaR})$ $\triangleright$ VaR along the path
@@
23:     $\textnormal{{VaR}}[v,\bar{\alpha}]\leftarrow\texttt{pathVaR}$ $\triangleright$ Update best estimate
24:     $\texttt{best\_paths}[v,\bar{\alpha}]\leftarrow\texttt{append}(\texttt{best\_paths}[v^{\prime},\alpha^{\prime}],v)$
25:     $\texttt{best\_samples}[v,\bar{\alpha}]\leftarrow\{y_{1},\dots,y_{n}\}$
@@
29:return $\textnormal{{VaR}}[t,\alpha],\texttt{best\_paths}[t,\alpha]$ $\triangleright$ Path with minimum VaR estimate
Towards optimizing the RMAG objective over an agent graph $G$, we remark that a simple procedure to find the optimal path is to estimate the $\textnormal{VaR}_{\alpha}$ along each path and pick the path that minimizes it. We can estimate the $\textnormal{VaR}_{\alpha}$ by computing the empirical $(1-\alpha)$-quantile on samples drawn from the loss along a path $L_{p}(Z_{p})$. While this is asymptotically optimal, it can be inefficient when the graph has an exponential number of paths as a function of the number of vertices.
To circumvent this, we present an efficient algorithm that only scales polynomially in the number of vertices in the agent graph and is also asymptotically near-optimal under certain assumptions. BucketedVaR is a dynamic programming algorithm that processes the vertices in the agent graph in the topological order and incrementally builds a near-optimal path.
To understand how it works, we first present how we can estimate the value-at-risk of the loss distribution of a single path incrementally. To this end, consider a path in an agent graph $p=v_{1}\xrightarrow{e_{1}}\dots\xrightarrow{e_{m}}v_{m+1}$ with $v_{1}$ being the source vertex and $v_{m+1}$ the target vertex. Let $(Z_{1},\dots,Z_{m})\sim\mathcal{D}(T_{e_{1}}\times\cdots\times T_{e_{m}})$ be a sequence of random variables representing the sequences of traces of the agents along the path. Then define $R_{1},\dots,R_{m}$ to be real-valued random variables representing the losses incurred by the agents along each edge. Formally, we can write $R_{1}\coloneqq L_{e_{1}}(Z_{1}),\dots,R_{m}\coloneqq L_{e_{m}}(Z_{m})$. If $L_{p}$ is the composed loss function along the path as defined in Equation˜2, we can rewrite $L_{p}(Z)$ as $\max(R_{1},\dots,R_{m})$. A key observation that enables the incremental estimation of $\textnormal{VaR}_{\alpha}[L_{p}(Z)]$ (or equivalently, the $(1-\alpha)$-quantile of $L_{p}(Z)$) is that we can estimate $\textnormal{VaR}_{\alpha_{i}}[R_{i}]$ of each loss variable individually with $\alpha_{i}$ chosen such that $\sum\alpha_{i}=\alpha$. Following this, we can apply the union bound to obtain
@@
This bound is justified in the proof of Theorem˜1. This method allows us to estimate the value-at-risk of each edge loss variable independently and then combine the estimates at the end by computing their maximum. However, we need to optimize the allocation of the risk budgets $\alpha_{i}$ along each edge. Intuitively, allocating a higher risk budget to higher edge losses would result in a lower value-at-risk for that edge since it would correspond to a lower quantile.
BucketedVaR optimizes this by searching for risk budget allocations in a discretized set of buckets $B=\{0,\frac{\alpha}{d},2\frac{\alpha}{d},\dots,\alpha\}$ for a given discretization factor $d\in\mathbb{N}_{>0}$. At each vertex $v$ and corresponding to each bucket $\bar{\alpha}\in B$, the algorithm inductively builds an optimal path that minimizes the upper bound on the value-at-risk along the partial path at a risk budget of $\bar{\alpha}$. It achieves this by considering each predecessor $v^{\prime}$ of $v$ and each allocation of budget $\alpha^{\prime}\leq\bar{\alpha}$ until $v^{\prime}$, for which we have inductively constructed the optimal path, and then estimates the value-at-risk by allocating a budget of $\bar{\alpha}-\alpha^{\prime}$ to the loss along the edge $v^{\prime}\to v$. The optimal path for $(v,\bar{\alpha})$ is then obtained by extending the optimal path to $(v^{\prime},\alpha^{\prime})$ that minimizes the value-at-risk estimate. A detailed proof of soundness and run-time analysis of the algorithm can be found in the proof of Theorem˜1
@@
[…]
@@
[…]
@@
Observe that since $\gamma\to 0$ as $n\to\infty$, the theorem tells us that the estimated value-at-risk $q$ is at least as large as the true value-at-risk of the loss distribution of the path $p$, which corresponds to $\textnormal{quantile}(L_{p}(Z_{p}),1-\alpha)$. We can show this using a union bound argument for the budget allocations found by the algorithm along with the fact that the empirical CDF converges uniformly to the true CDF by the DKW inequality [8, 23]. The time complexity is easily deduced by looking at the pseudocode of BucketedVaR. See Appendix˜A for the complete proof. ∎
Theorem˜1 only justifies the validity of the union bounding method for estimating the value-at-risk but we have not discussed the optimality of BucketedVaR. Here, we can only hope to achieve tight estimates of the value-at-risk when the loss variables along every path are independent or are only loosely correlated. Under these independence assumptions, Theorem˜2 tells us that asymptotically, the quantile estimated by the BucketedVaR algorithm is at most the $(1-\alpha+\alpha^{2}/2)$-quantile of the losses along the optimal path. In other words, the path found by our algorithm is suboptimal by an additive factor of $\alpha^{2}/2$ with respect to the quantiles of the optimal path loss.
@@
This is again a consequence of the uniform convergence of the empirical CDF along with the fact that the independence of the loss variables allows us to deduce an upper bound. See Appendix˜A for the complete proof. ∎
@@ 4.1 Benchmarks
We implement the BucketedVaR algorithm for risk minimization and evaluate its performance on a suite of discrete and continuous control benchmarks that require composing agents trained with reinforcement learning (RL) to complete long-horizon objectives. In all the benchmarks that we consider, the agent graphs $G=(V,E,X,T,F,L,s,t,\mathcal{D}_{s})$ share the following structure: the domain associated with each vertex $X_{v}$ is the state space of the control environment $\Sigma$ and the agents along each edge are control policies $\pi_{e}:\Sigma\to\mathfrak{D}(\Sigma^{*}\times\Sigma)$ that take an initial state and produce a distribution over trajectories and final states. These final states then become the initial states for the next policy. Lastly, $\mathcal{D}_{s}$ represents the distribution over the initial states of the environment.
@@
Maximizing safety during navigation and manipulation. We test our algorithm on the 16-Rooms and Fetch environments from Jothimurugan et al. [18] who present a compositional RL framework to complete long-horizon objectives in continuous environments. An objective is specified as a task graph which is the same as an agent graph. In the task graph, every edge represents a reach-avoid subtask: reaching a target region while avoiding a set of dangerous states. Each reach-avoid subtask specifies a real-valued reward function that takes in a trajectory of execution of the policy and evaluates to at least zero if the trajectory successfully completes the subtask. Intuitively, the reward captures a distance metric between trajectories and reach-avoid sets. The loss function in the agent graph is defined as the negative of the reward function.
@@
The 16-Rooms environment consists of 16 rooms arranged in a $4\times 4$ grid with doors between adjacent rooms. RL policies are trained to control a point mass to complete reach-avoid tasks. The agent graph has 13 vertices and 16 paths of length 8 each. The second Fetch environment [26] is taken from Gymnasium-Robotics and involves controlling a 7-DoF robotic arm with a gripper attached at the end to pick up and move objects. We consider the task graph for the PickAndPlaceChoice objective that directs the gripper to move near the object, grip it, pick it up and move along one of two trajectories. The graph has 7 vertices and 2 paths of length 5. Additionally, we adapt the Rooms environment to implement a version of the drone navigation task from Example 1.
Figure 2: Top-view map with agent graph of the BoxRelay environment.
@@
Minimizing resource consumption. We use the Miniworld [5] framework to implement the BoxRelay environment of which the top-view map is shown in Figure˜2. Observations in this environment are given as first-person view RGB images (see Section˜B.4). The objective is to move the player (red triangle) to the first box (marked 1) and pick it up. The player can then choose to go to either of the two boxes in the two middle rooms (3 or 7), place the initial box, pick up the next box, and finally drop it at the location marked in the right room with a star. Furthermore, after picking up the box at 7, the player can choose to exit through one of the two exits 8 or 9. We add the constraint that the player has limited battery power but gets recharged after dropping the first box and before picking up the second box. So we would like to minimize the amount of time that either box is held by the player. Importantly, since the initial positions of the player, the boxes, and the final target are all chosen uniformly at random within each room, this is a non-trivial optimization problem.
@@
We can model the BoxRelay task as an agent graph as follows: the graph is as shown in Figure˜2 with the numbers being the vertices. The edges represent the agents which are RL policies trained to navigate the player to pick up and place objects at the desired locations. The loss functions evaluate to the number of time steps that a box is carried before being put back down and thus quantifies the resource consumption. Here, the RMAG objective corresponds to minimizing the maximum time that any box is carried by the player.
4.2 Experimental results
@@
Equipped with these benchmark environments, we conduct three sets of experiments to evaluate the effectiveness of the approximation obtained from the BucketedVaR algorithm. We measure the accuracy of the approximation by estimating the empirical quantile on $10^{4}$ fresh samples of losses along the optimal path. Specifically, if $c\in\mathbb{R}$ is the approximate $\textnormal{VaR}_{\alpha}$ returned by the BucketedVaR algorithm, we draw samples of losses along the optimal path and compute the fraction of these which are below $c$ giving us the empirical quantile of $c$. We say that the approximation is tight if the empirical quantile is close to the desired quantile $1-\alpha$.
Figure 3: Absolute quantile difference in percentage between the quantile computed by the BucketedVaR algorithm and the desired quantile. The empirical quantiles are computed on a fresh set of $10^{4}$ samples.
Experiment 1: Comparison with optimal baseline algorithm. We define a baseline algorithm that explicitly esimates $\textnormal{VaR}_{\alpha}$ at the desired risk level $\alpha$ along each path of the agent graph. While this algorithm is asymptotically optimal, it is inefficient as discussed in the previous section. The number of samples for both algorithms is set to $10^{4}$ and the number of buckets is chosen based on the number of agents to allow sufficient granularity in the allocation of the risk budget. For example, in the 16-Rooms benchmark since there are 8 agents along each path, we choose a higher number of buckets.
We find that BucketedVaR succeeds in finding the same optimal path as the baseline algorithm across all the benchmarks along with tight estimates of $\textnormal{VaR}_{\alpha}$. Figure˜3 shows the absolute difference in percentage points of the quantile estimates obtained by the BucketedVaR algorithm and the desired quantiles. We remark that the difference is never more than a couple percentage points across all benchmarks and risk levels. It also finds non-trivial allocations of the risk budget for the agents: in the 16-Rooms benchmark, the budget allocations for the 8 agents along the path to estimate $\textnormal{VaR}_{0.1}$ are $16\bar{\alpha},0\bar{\alpha},10\bar{\alpha},23\bar{\alpha},19\bar{\alpha},11\bar{\alpha},7\bar{\alpha}$, and $14\bar{\alpha}$ for $\bar{\alpha}=0.1/100$.
Experiment 2: Increasing sample size and number of buckets. Increasing the sample size improves the approximation because we have more accurate quantile estimates and similarly a higher number of buckets results in more granular risk budget allocations. This is confirmed by the two experiments we run: in the first one, plotted in Figure˜4(a), we maintain the number of buckets at 100 and vary the number of samples from 500 to $10^{4}$ with $\alpha=0.1$. The empirical quantile quickly stabilizes around 0.91 across benchmarks. The second experiment maintains the sample size at $10^{4}$ and varies the number of buckets from 5 to 100 with $\alpha=0.1$. Figure˜4(b) again shows a similar trend. We note that for the 16-Rooms benchmark, a low number of buckets gives bad VaR estimates since the number of agents along each path is higher.
Experiment 3: Scaling number of agents. From our theoretical result, we expect BucketedVaR to continue producing tight estimations of the VaR when we increase the number of agents along the path, provided that the condition on the independence of losses holds and the sample size is large enough. We test this hypothesis by simulating a long path of agents using the 16-Rooms benchmark. We consider the optimal path in this benchmark which has a sequence of 8 agents and construct a new agent graph with $8k$ agents for $k\in\{1,2,3,4,5\}$. We can sample a sequence of trajectories from this path by taking $k$ trajectories of the original 8 agents. The losses are then computed using the original loss functions. Since these are independently drawn trajectories, the assumption of independence holds. We estimate the $\textnormal{VaR}_{\alpha}$ of this path of agents using BucketedVaR for $10^{4}$ samples, $\alpha=0.1$, and $100$ buckets. The results are plotted in Figure˜4(c) in which we observe that the approximation continues to remain tight with an increasing number of agents.
Takeaways. Through our benchmarks, we illustrated that we can use our framework to model risk minimization associated with agent compositions with respect to safety and resource consumption requirements. Then we showed that the BucketedVaR algorithm produces estimates of the value-at-risk that are very close to the desired level of $1-\alpha$. Furthermore, we confirmed our theoretical result that the approximation improves with growing sample size and granularity of the risk budget division given by the number of buckets.
@@
Figure 4: Empirical quantiles of the $\textnormal{VaR}_{0.1}$ estimates computed by BucketedVaR with varying parameters. $95\%$ Clopper-Pearson CIs for the empirical quantiles computed on $10^{4}$ samples are also plotted.